OpenStudy (anonymous):

Find all of the zeros of the function: f(x)=-3(5x-6)(2x+7)(x-4)

OpenStudy (anonymous):

you have three terms : (5x-6), (2x+7), and (x-4). Set each of these terms to zero and solve for x

OpenStudy (anonymous):

Ok, let me try that

OpenStudy (anonymous):

post what you get

OpenStudy (anonymous):

Ok, I got, x=(6/5) x=-(7/2) AND x=4

OpenStudy (anonymous):

excellent

OpenStudy (anonymous):

Ok, now what?

OpenStudy (anonymous):

thats it, those are the zeros of that function

OpenStudy (anonymous):

What happened to the -3 though?

OpenStudy (anonymous):

what happens when you set -3 to be zero

OpenStudy (anonymous):

there is no x to solve for

OpenStudy (anonymous):

So the -3 means nothing?

OpenStudy (anonymous):

the -3 does not contribute a zero

OpenStudy (anonymous):

It makes sense, but would you possibly times it through the parentheses?

OpenStudy (anonymous):

So 3 zeros?

OpenStudy (anonymous):

no, the -3 does not contribute anyhting

OpenStudy (anonymous):

koolkimberly, why dont you multiply into one of the terms and tell me what you get when you set that term to zero

OpenStudy (anonymous):

Are you a teacher?

OpenStudy (anonymous):

no, just a student like your self

OpenStudy (anonymous):

what did you get koolkimberly?

OpenStudy (anonymous):

Wow......I got the same answer, 6/5! Cool! Thanks LagrangeSon!!!